### v2

# SPIN-ODE: Stiff Physics-Informed Neural ODE for Chemical Reaction Rate Estimation

Abstract
Estimating rate coefficients from complex chemical reactions is essential for advancing detailed chemistry. However, the stiffness inherent in real-world atmospheric chemistry systems poses severe challenges, leading to training instability and poor convergence that hinder effective rate coefficient estimation using learning-based approaches. To address this, we propose a Stiff Physics-Informed Neural ODE framework (SPIN-ODE) for chemical reaction modelling. Our method introduces a three-stage optimisation process: first, a latent neural ODE learns the continuous and differentiable trajectory between chemical concentrations and their time derivatives; second, an explicit Chemical Reaction Neural Network (CRNN) extracts the underlying rate coefficients based on the learned dynamics; and third, fine-tune CRNN using a neural ODE solver to further improve rate coefficient estimation. Extensive experiments on both synthetic and newly proposed real-world datasets validate the effectiveness and robustness of our approach. As the first work on stiff Neural ODEs for chemical rate coefficient discovery, our study opens promising directions for integrating neural networks with detailed chemistry.
I Introduction
Neural Ordinary Differential Equations (Neural ODEs) have emerged as a powerful approach to bridging mathematical dynamical systems with neural networks. Their key advantage lies in representing system dynamics through continuous-depth neural networks. Neural ODEs have shown significant promise in the domains of physics and chemistry [1, 2, 3, 4]. However, despite these advances, their application to detailed chemical modelling remains limited in the computer science community, particularly in tasks such as reaction rate coefficient estimation and the discovery of novel chemical pathways [5].
Figure 1: The proposed SPIN-ODE fits concentration trajectories (a) to infer reaction rate coefficients (b), converging from random initialisation to the true data (light colour $\rightarrow$ dark colour).
Understanding and predicting chemical reaction rate coefficients is a cornerstone of detailed chemical modelling. Unlike implicit modelling approaches [6, 7, 8, 9, 3] that focus on approximating system-level outputs, explicitly learning reaction rate coefficients demands a nuanced understanding of chemical kinetics. This task is particularly difficult in computer science, where knowledge transfer from domain-specific numerical chemistry to neural networks is still in its infancy. The fundamental obstacle lies in identifying and encoding the essential factors, such as temperature dependencies and chemical reaction paradigms, into a learnable and generalisable model architecture.
Moreover, the scarcity of real experimental data further complicates this endeavour. Most available measurements come from highly controlled environments like flow-tube [10] or smog-chamber [11] experiments, which are costly, time-consuming, and often prone to significant measurement uncertainties. As a result, purely data-driven learning approaches suffer from limited generalisation and a lack of physical grounding. Furthermore, chemical reaction systems are often inherently stiff, characterised by vastly different reaction timescales, which significantly complicates numerical integration and destabilises neural network training. To address this, our method incorporates domain-informed knowledge derived from detailed numerical chemical models, allowing the neural network to learn subject to physical constraints and chemical principles even in data-sparse and numerically stiff regimes.
In this work, we propose a novel framework that utilises the strengths of Neural ODEs to estimate reaction rate coefficients. As illustrated in Figure 1, our keen interest is to discover the correlations between chemical concentration trajectories and rate coefficients, so that the optimal rate coefficients can be found for ODE fitting. Rather than using neural networks as an implicit ODE system, our goal is to discover the hidden chemical parameters for detailed chemistry. We envision this integration of AI and detailed chemistry for scientific discovery, potentially uncovering new patterns, simplifying mechanisms, and offering interpretable insights into complex reaction systems. To summarise:
  • We propose an explicit stiff physics-informed neural ODE framework (SPIN-ODE) for detailed chemistry. This framework efficiently models ODE trajectories while simultaneously extracting interpretable reaction rate coefficients.
  • We introduce a three-stage optimisation strategy specifically designed to handle stiffness and enhance training stability. Our novel physical loss functions and trajectory resampling techniques significantly improve ODE fitting and rate coefficient estimation.
  • Compared to existing approaches that primarily calibrate rate coefficients of stiff reactions based on initial close guesses, our method is the first work to automatically find optimal rate coefficients in the absence of prior knowledge.
II Related Work
Our work is on detailed chemical kinetics via neural networks, which is closely related to neural networks and ODEs, and their applications in chemical modelling. Let us revisit those topics in detail.
II-A Numerical and Neural ODEs
Numerical methods for solving ordinary differential equations (ODEs) have been extensively studied, with classical approaches such as Euler’s method, Runge-Kutta schemes, and multistep methods [12] forming the foundation of the field. These techniques aim to approximate solutions to initial value problems with varying degrees of accuracy, stability, and computational efficiency. Over the decades, significant advances have been made in adaptive step-size control, stiff equation solvers, and error estimation strategies. Recent developments have also explored the intersection of numerical analysis and machine learning, where neural networks and physics-informed approaches offer data-driven alternatives for solving ODEs. Neural ODEs were first introduced in [13], where backpropagation can be used to train an ODE model. Later on, many works [14, 15, 16, 17, 18, 19] followed up on this direction to further analyse and improve the performance. For example, [15] analyses the theory behind the infinite-depth neural ODEs and proposes adaptive-depth and data-controlled neural ODE paradigms. [18] demonstrates that a simple augmentation of the latent ODE space can achieve lower losses with few parameters and is more stable to train. To overcome the slow training process, [14] proposes a regularised neural ODE (RNODE) to learn well-behaved dynamics.  [20] proposes Fourier analysis to ODEs for temporal estimation and potentially high-order spatial gradients calculation. Later on, Neural ODEs are also applied to domain-specific applications, like time-series processing [16], latent space image processing [17], glucose dynamics [21] and robotics control [22].
Meanwhile, a significant challenge to the current Neural ODEs is the stiff ODE system. [23] was the first to observe that explicit methods fail for certain chemical kinetic systems. The ”stiff“ chemical reactions can happen so fast that small errors can be magnified if the equations are integrated. Kim et al., [24] study on two classical stiff chemical systems using their proposed scaling and adjoint solution to mitigate stiffness. Fronk et al. present two consecutive studies [25, 26] to demonstrate both explicit and implicit solution to neural stiff ODE training. Physics-informed neural network (PINN) [27] is another solution by integrating more physical constraints to tackle the high-stiff regime. Ji et al. [28] apply steady state analysis in the physics-informed loss to reduce the stiffness of the reaction system. Seiler et al. [29] propose a method based on transfer learning and design a multi-head PINN network to divide and conquer the ODE system into low-stiff and high-stiff regions. Sulzer et al. [30] propose an autoencoder to solve stiff ODE in the latent space. The first- and second-order gradients are used as physical losses for optimisation.
II-B Chemical Neural Networks
The rise of deep learning in the early 2010s has significantly expanded the scope of scientific discovery processes. Neural networks for chemical modelling have recently attracted significant research interest, like weather forecasting [1, 31, 32], air quality estimation [8, 6, 33, 7], and others. Chemical kinetics can be considered as a complex ODE system and researchers have been studying both implicit and explicit approaches for efficient chemical modelling. For example, Owoyele and Pal [3] propose to utilise neural ODE to simulate complete chemical kinetics implicitly for fast computation. DeepONet [2] integrates the recently proposed Neural Operator [34] into the neural network to exploit the dynamical structure of chemical reactions. Liu et al. [4] further improve it by proposing ChemODE, an attention based neural operator for fast and accurate chemical modelling, which can successfully estimate over hundreds of chemical reactions. Kolmogorov-Arnold Networks (KAN) were proposed recently [35] as an alternative to multi-layer perceptrons for general neural network applications. It attracts a lot of attention for chemical modelling [36]. The advantage of using KAN is that it provides fast neural scaling and ”white-box“ interpretability and generalisability. ChemKAN [9] was later proposed to learn hydrogen combustion chemistry modelling, achieving a twofold speed-up over detailed numerical chemistry.
In contrast to implicit models that approximate the overall chemical system behaviour, explicit chemical neural networks aim to directly learn or represent the fine-grained chemical mechanisms themselves, such as reaction rates, activation energies, or reaction pathways. This explicit modelling is crucial for advancing scientific understanding, especially in domains where interpretability and physical consistency are as important as predictive performance. Despite its importance, relatively few studies have focused on using neural networks to capture detailed chemical kinetics in an interpretable and physically faithful manner. This gap is largely due to the complexity of chemical reaction networks and the challenge of integrating domain knowledge, such as thermodynamic constraints and reaction mechanisms, into machine learning models. Some notable efforts include ChemTab [5], which tabulates reaction rates using neural networks for combustion modelling, and ReactionNet [37], which predicts reaction rates by incorporating reaction templates and thermochemical data. Ji et al. [38] autonomously discover reaction pathways by regression on stoichiometric and rate coefficients. However, it is only used to calibrate the rate coefficients initialised with a small-scale uncertainty given fixed reaction pathways [39]. In this work, we argue that explicit neural representations of chemical reactions offer a promising direction for coupling accuracy, interpretability, and computational speed in scientific modelling, especially when integrated with recent developments in symbolic neural networks or physics-informed learning frameworks.
III Approach
Figure 2: Our proposed SPIN-ODE framework for stiff reaction rate coefficient estimation. It includes three steps: 1) latent neural ODE fitting by optimising predicted ODE trajectories, 2) resampling data from step 1 and pre-training physics-based CRNN by optimising predicted time derivatives, and 3) fine-tuning CRNN with the ODE solver by optimising predicted concentration trajectories. The detailed structure of the CRNN is shown in the upper right corner.
III-A Preliminaries
1. Chemical kinetics and stiff ODEs. We illustrate the chemical kinetics with the following example Robertson [40] reaction system:
[…]
Without loss of generality, any of these reactions can be represented in a unified form as:
$s_{i,a}^{f}\textrm{A}+s_{i,b}^{f}\textrm{B}+s_{i,c}^{f}\textrm{C}\rightarrow s_{i,a}^{r}\textrm{A}+s_{i,b}^{r}\textrm{B}+s_{i,c}^{r}\textrm{C},$ (2)
in which $s_{i,j}^{f}$ and $s_{i,j}^{r}$ are the forward and reverse stoichiometric coefficients for species $j$ in reaction $i$, indicating molecules consumed and produced per reaction. The reaction rates are determined by a power-law expression involving the rate coefficient $k_{i}$ and reactant concentrations:
$r_{i}=k_{i}[A]^{s_{i,a}^{f}}[B]^{s_{i,b}^{f}}[C]^{s_{i,c}^{f}}.$ (3)
Here we use []delimited-[][$\cdot$][ ⋅ ] to represent chemical concentrations. Reaction rate coefficient $k_{i}$ follows the Arrhenius relationship and is dependent on the temperature and molecular energy barriers, and is often determined experimentally. The overall rate of change for species concentration $Y_{j}$ is the sum of the consumption and production rates from all reactions:
$\frac{d[Y_{j}]}{dt}=\sum_{i}-s_{i,j}^{f}r_{i}+s_{i,j}^{r}r_{i}.$ (4)
Therefore, the governing equations for the Robertson reaction system, shown in Eq (1), constitute a system of non-linear ODEs:
$\frac{d[A]}{dt}$ $=-k_{1}[A]+k_{3}[B][C],$ (5)
$\frac{d[B]}{dt}$ $=k_{1}[A]-k_{2}[B]^{2}-k_{3}[B][C],$
$\frac{d[C]}{dt}$ $=k_{2}[B]^{2}.$
When reaction timescales vary significantly, solving the ODE with ordinary numerical methods becomes unstable, and the system is identified as numerically stiff. The stiffness of a system is often characterised by the stiffness ratio, defined as the disparity between the maximum and minimum eigenvalues $\lambda$ of the system Jacobian:
Stiffness $\text{Stiffness ratio}=\frac{\max|Re(\lambda)|}{\min|Re(\lambda)|}.$ (6)
High stiffness ratios are typical in atmospheric chemistry. Notably, in cases like the classic stiff Robertson problem, zero eigenvalues result in the stiffness ratio being undefined. While numerical stiff ODE solvers are extensively studied, their robust integration into machine learning frameworks remains limited.
2. Neural ODEs. Similar to chemical reactions, many systems across science and engineering domains can be represented by ODEs:
$\frac{dy(t)}{dt}=f(y(t),t),$ (7)
where $y(t)$ is the system state variable at time $t$, and the function $f$ governs the system dynamic. Neural Ordinary Differential Equations (Neural ODEs) approximate unknown system dynamics $f$ using neural networks parametrised by $\theta$:
$\frac{dy(t)}{dt}=NN_{\theta}(y(t),t).$ (8)
Since observational data typically includes only discrete-time state variables $y(t)$, Neural ODEs integrate the approximated dynamics to predict system states at subsequent times:
[…]
$=ODESolver(NN_{\theta},y(t_{0}),t_{0},t_{1}).$
The loss function is defined as the distance, e.g., mean squared error (MSE), between model predictions and observed state variables. With the differentiable $ODESolver$, the gradient of the loss with respect to the network parameters $\theta$ can be calculated and leveraged to optimise the network using gradient-based algorithms. Chen et al. propose a widely-used approach for memory-efficient gradient calculation by solving the backward-in-time [13]. However, neural ODE training is extremely vulnerable to stiffness, where minor errors in the forward pass are rapidly amplified in the backward pass [24]. In this paper, we are specifically interested in solving stiff ODEs in atmospheric chemistry for rate coefficient estimation. Existing stiff neural ODE solutions, like collocation based ODE training [41, 42] and temporal regularization based ODE training [19] fail to fit ODE trajectories, and cannot extract the rate coefficients.
III-B Proposed method
We propose a novel neural network approach, SPIN-ODE, for stiff chemical modelling, which not only learns ODE fitting, but also explicitly extracts the rate coefficients for detailed chemistry. As shown in Figure 2, the whole network consists of three modules: MLPs, ODE solver, and Chemical Reaction Neural Network (CRNN). The overall process includes three steps: 1) latent ODE fitting, 2) resampling ODE trajectories for physics-based CRNN pre-training, and 3) fine-tuning CRNN with the ODE solver. The first step is to train MLPs to implicitly estimate ODE trajectories given input initial chemical concentrations. The second step is to resample from learned ODE trajectories to pre-train the CRNN for rate coefficient approximation. The final step is to refine rate coefficients by combining CRNN and an ODE solver for fine-tuning. Let us introduce them in detail.
SPIN-ODE step 1: Latent Neural ODE fitting In the first step, we fit the trajectory with a latent neural ODE network. We combine an MLP module and an ODE solver to implicitly learn the neural ODE curve fitting as Eq 9. Once it is trained, the MLP approximates the correlations between $y$ and $dy/dt$. The key component is the design of the denormalisation layer, which is crucial for ODE fitting in stiff chemistry. The motivation comes from [24], where the stiffness of the concentration roughly aligns with their time derivatives, such that they need to scale up to the same range for stable optimisation. Specifically, we describe the process as follows:
$\frac{dy(t)}{dt}=NN_{\theta}(\frac{y(t)-y_{\text{min}}}{y_{\text{max}}-y_{\text{min}}},t)\frac{y_{\text{max}}-y_{\text{min}}}{t_{\text{scale}}},$ (10)
where $y_{\text{max}}$ and $y_{\text{min}}$ are the species-wise maximum and minimum concentrations used in the normalisation layer, and $t_{\text{scale}}$ is the time span of the entire measurements. The ODE solver integrates the output time derivative to get a trajectory as Eq (9), and the loss function is correspondingly normalised by the scale of concentrations as:
$\mathcal{L}_{y}=\text{MSE}(\frac{y(t)^{\text{model}}}{y_{\text{max}}-y_{\text{min}}},\frac{y(t)^{\text{observe}}}{y_{\text{max}}-y_{\text{min}}}).$ (11)
In addition to the loss on predicted state variables, we add two physical loss terms on the velocity and acceleration of the concentration change, enabling the learned dynamics to be further constrained by the physical laws. Specifically, we compute the 1st-order and 2nd-order derivatives losses ($\mathcal{L}_{\dot{y}}$ and $\mathcal{L}_{\ddot{y}}$) with respect to time, then apply the same scaling technique as in Equation (11), and the final loss function is:
$\mathcal{L}=\mathcal{L}_{y}+\mathcal{L}_{\dot{y}}+\mathcal{L}_{\ddot{y}}.$ (12)
SPIN-ODE step 2: Resampling and pre-training CRNN. While non-linear system dynamics can often be accurately modelled using simple deep neural networks, i.e., MLPs, the parameters learned by such black-box models typically lack interpretability. In contrast, physics-based neural ODEs encode parameters and mathematical operators derived directly from physical equations [38, 43]. Following [38, 39], we leverage the Chemical Reaction Neural Network (CRNN) (see Figure 2), a physics-based neural network tailored for discovering chemical reaction parameters. CRNN expresses the reaction rate law in Eq (2) as a linear combination of the logarithm of the reaction rate coefficient and the logarithms of species concentrations, followed by an exponential transformation. Each network node represents an individual reaction, and nodes can be stacked to form a comprehensive model of a reaction system. The overall time derivative of species concentrations $d\mathbf{Y}/dt$, is obtained by linearly transforming the reaction rate vector $\mathbf{R}$, through the net stoichiometric matrix $\mathbf{S}$, defined as $s_{i,j}=-s_{i,j}^{f}+s_{i,j}^{r}$:
$r_{i}=\exp\left(\ln(k_{i})+s_{i,a}^{f}\ln[A]+s_{i,b}^{f}\ln[B]+s_{i,c}^{f}\ln[C]\right).$ (13)
$\frac{d\mathbf{Y}}{dt}=\mathbf{S}\cdot\mathbf{R}.$ (14)
To facilitate CRNN training, we first perform dense sampling of data points from the pre-trained MLPs in step 1, and then use the resulting “augmented points” for supervised learning. Specifically, when the latent neural ODE fits the concentration trajectory ${{y}(t_{i})}_{i=1}^{M}$ with original $M$ points, we increase the time resolution and integrate an interpolated concentration trajectory ${\hat{y}(t_{i})}_{i=1}^{N}$ with $N>M$. We then apply the Central Difference Formula (CDF) to estimate the corresponding time derivative $\dot{\hat{y}}(t)$ at state $\hat{y}(t)$:
$\dot{\hat{y}}(t_{i})\approx\frac{\hat{y}(t_{i+1})-\hat{y}(t_{i-1})}{t_{i+1}-t_{i-1}},\quad\text{for }1<i<N,$ (15)
with forward and backward difference schemes applied at the endpoints $i=1$ and $i=N$ respectively:
$\dot{\hat{y}}(t_{1})$ $\approx\frac{\hat{y}(t_{2})-\hat{y}(t_{1})}{t_{2}-t_{1}},$ (16)
$\dot{\hat{y}}(t_{N})$ $\approx\frac{\hat{y}(t_{N})-\hat{y}(t_{N-1})}{t_{N}-t_{N-1}}.$ (17)
Using this augmented dataset of $\left(\hat{y}(t_{i}),\dot{\hat{y}}(t_{i})\right)$, we can formulate a supervised learning task using the same scaled MSE loss as Eq 11.
The reason behind such ”first-regression-then-estimation“ design is twofold: 1) the physical interpretability of network parameters depends on their alignment with the corresponding physical equations, thus precluding operations that violate physical laws; and 2) for non-linear chemical kinetics, we apply linear normalisation directly to input concentrations and denormalisation in step 1 for stable optimisation. It induces non-linear distortions of reaction rates, causing deviations of learned parameters from their true physical meaning.
SPIN-ODE step 3: Fine-tuning CRNN with ODE solver. After the CRNN is pre-trained, we resume the training by combining the CRNN and the ODE solver to fine-tune the rate coefficients. The loss function is as the same as that in Eq 12. The initial learned rate coefficients during step 2 avoid the extreme error growth and stabilise the ODE integration. We finally extract the learned rate coefficients $k_{i}$ in the physics-based CRNN as the final estimation.
IV Experiments
Dataset Robertson POLLU AOXID
Task ODE fitting Rate coeff. estimate ODE fitting Rate coeff. estimate ODE fitting Rate coeff. estimate
Ji et al. [38] $2.2\times 10^{-2}$ 1.93 $9.3\times 10^{19}$ 9.25 * *
SPIN-ODE-s1 $2.8\times 10^{-3}$ $1.1\times 10^{-5}$ $1.4\times 10^{-5}$
SPIN-ODE-s2 $3.4\times 10^{-3}$ 0.69 $1.9\times 10^{-2}$ 1.42 $8.9\times 10^{-5}$ 0.15
SPIN-ODE-s3 $6.3\times 10^{-6}$ $7.4\times 10^{-3}$ $8.6\times 10^{-6}$ 0.46 $4.5\times 10^{-6}$ 0.10
TABLE I: Overall comparisons between our proposed three-step SPIN-ODE approach and Ji et al. [38] model. We show the scaled MSE loss of the ODE trajectory fitting, as well as the log-MAE loss on rate coefficient estimation. - means the network involved in the current step does not provide estimation, while * means runtime error.
IV-A Experimental setups
Our method is first validated using the Robertson problem, a classical benchmark comprising three chemical species and three highly stiff reactions as presented in Eq (1). The governing equations for the Robertson problem are explicitly defined in Eq (5), with ground-truth reaction rate coefficients as $[0.04,3\times 10^{7},1\times 10^{4}]$.
Real-world atmospheric chemistry systems typically involve significantly more complex dynamics, exhibiting substantial stiffness across numerous reactions. Consequently, we further evaluate our method on the POLLU problem [44], a widely utilised atmospheric chemistry model for stiff ODE benchmarks. This system includes 25 reactions among 20 chemical species, with reaction rate coefficients spanning from $10^{-3}$ to $10^{12}$. In addition, we introduce an autoxidation reaction (AOXID) scheme involving 44 species and 49 reactions. It is a simplified system representing realistic autoxidation processes in the atmosphere characterized by intricate radical chain mechanisms [45]. The reaction rate coefficients in this system vary dramatically, ranging from approximately $10^{-14}$ to $10^{6}$. Comprehensive details for the POLLU and AOXID reaction schemes are provided in the supplementary materials.
Datasets for the Robertson and POLLU problems are numerically generated using a stiff-aware Backward Differentiation Formula (BDF) solver available in the SciPy library. The AOXID dataset is generated using the Rosenbrock solver within the well-known Kinetic PreProcessor (KPP) [46]. For the Robertson system, initial concentrations are set as $[y_{1},y_{2},y_{3}]=[1,0,0]$, with 50 data points sampled logarithmically spaced between $10^{-5}$ and $10^{5}$ seconds. For the POLLU and AOXID datasets, initial conditions detailed in the supplementary material are used, and 100 data points are uniformly sampled across the time span of 0.1 and 100 seconds, respectively. To mitigate numerical instabilities arising from prolonged integration and accumulation of errors, a sliding window approach of size 20 and stride of 10 is adopted during the training for POLLU and AOXID.
The proposed SPIN-ODE leverages the stiff-aware differentiable ODE solver Kvaerno3 from the diffrax library [47]. The MLP module is based on a 3-layer Multilayer Perceptron with 128 nodes per hidden layer and GELU activation. Experiments are conducted using an NVIDIA A100 GPU with 40 GB of memory. Network parameters are optimised using the Adam algorithm, initialised with a learning rate of 0.001, which is annealed when training loss does not decrease for 10% of the total epochs. The code is available at https://github.com/pvvq/SPIN-ODE.
IV-B Comparison with state of the art
Our overall comparisons include both ODE trajectory fitting and rate coefficient estimation. Given the wide-ranging magnitudes of the chemical concentration and reaction rate coefficients, we evaluate the accuracy of model estimations using the mean absolute error (MAE) on a logarithmic scale. To the best of our knowledge, we are the first to study stiff chemical ODE for rate coefficient estimation. There is only one related work, Ji et al. [38], which we reimplement for comparisons.
The quantitative results on ODE fitting and rate coefficients are reported in Table I. We show our results in every step (SPIN-ODE-s1, SPIN-ODE-s2, SPIN-ODE-s3). As shown in Table I, ours outperforms Ji et al. [38] in every step. It is interesting to see that Ji et al. fail to converge on the AOXID dataset, and large error on the POLLU dataset. On the other hand, our final results (SPIN-ODE-s3) achieve low errors in rate coefficient estimation and ODE fitting, specifically $7.4\times 10^{-3}$ for Robertson, $8.6\times 10^{-6}$ for POLLU, and $4.5\times 10^{-6}$ for AOXID. The promising results on the AOXID dataset indicate that our approach could be used in real-world atmospheric chemistry. Furthermore, by comparing SPIN-ODE-s3 with SPIN-ODE-s2, it demonstrates the benefit from the final refinement process. Note that in SPIN-ODE-s1, we use MLPs to implicitly learn ODE fitting so that it cannot estimate rate coefficients yet. Comparing SPIN-ODE-s1 and SPIN-ODE-s2, the latter increases the ODE fitting errors because it learns to search for good rate coefficients to fit the physics-based CRNN model. The CRNN has fewer free parameters than MLPs to minimize the overall ODE fitting errors. However, once the optimal values of $k$ is found, SPIN-ODE-s3 produces the lowest errors in the ODE fitting.
To better visualise the ODE trajectory fitting, we plot the trajectory predictions in Figure 3. From left to right, we have the results for Robertson, POLLU and AOXID. Note that we select the three most representative reactions from POLLU and AOXID for visualization. The complete reaction plots can be found in the supplementary material. Since the method by Ji et al. performs reliably only on the Robertson problem, we did not show its trajectories on POLLY and AOXID. For our approach, it fits the ground truth trajectories in all three cases.
Since our primary focus is on rate coefficient estimation, we also plot the scatter diagram to demonstrate all rate coefficients in Figure 4. We can see that the model predictions (blue cross) closely match with ground truth values (yellow circles). Note that for the POLLU problem, coefficients 15-19 correspond to oxidant cycling, whose rate coefficients have been intensively studied and are provided during training and thus excluded from the estimation target.
Figure 3: Concentration trajectory fitting compared with observation. From left to right are the results for Robertson, POLLU and AOXID problems. 3 representative concentration trajectories are selected from the POLLU and AOXID problems. Results for Robertson show a gradual approximation to ground truth with our incremental approach.
Figure 4: Estimated rate coefficients for Robertson, POLLU and AOXID, from left to right. The blue cross symbol represents the model prediction, and the yellow circle represents the ground truth value.
IV-C Ablation studies
We compare four strategies for our pre-training step to illustrate the proposed key components: 1) we directly apply the Central Difference Formula (CDF) to the original observation and obtain the time derivative without SPIN-ODE-s1 MLP fitting, we refer to it as Direct CDF, 2) after MLP fitting, we directly use MLP outputs to train the CRNN, without SPIN-ODE-s2 resampling and differentiating, hereafter referred to as MLP proxy. And 3) in SPIN-ODE-s2, we propose to resample from the learned trajectories and then apply CDF. We evaluate the pre-training result both with and without dense interpolation. To ensure fairness, all methods without interpolation are trained for proportionally longer epochs to match the number of collocation points.
ODE fitting Rate coeff. estimate
Direct CDF $3.4\times 10^{-4}$ 0.192
MLP-proxy $1.1\times 10^{-6}$ 0.167
SPIN-ODE-s2 w/o Interpolation $7.4\times 10^{-5}$ 0.173
SPIN-ODE-s2 w/ Interpolation $8.9\times 10^{-5}$ 0.145
TABLE II: Ablation studies. We report the quantitative results for our pre-training step on the AOXID dataset with or without key proposed components.
Table II presents the results on the AOXID dataset. The CDF method applied directly to sparse observations performs worst in both trajectory prediction and parameter estimation, reflecting its sensitivity to data resolution. While using the latent network output as a proxy for derivatives yields the best trajectory match with $\mathcal{L}_{y}=1.1\times 10^{-6}$, its higher $\mathcal{L}_{k}$ suggests that accurate reproduction of concentration trajectories does not necessarily imply reliable estimation of reaction coefficients. This discrepancy arises because neural ODE solvers perform integration over multiple internal steps between successive observation points, and the model receives no supervision at these intermediate time steps. Row 3 performs comparably to row 2, indicating no additional information from fitting the MLP without interpolation. Interpolation in trajectory space followed by differentiation can be interpreted as a smoothing operation in time-derivative space. The last row achieves the lowest rate coefficient error $\mathcal{L}_{k}=0.145$, indicating the most accurate recovery of system dynamics.
IV-D Robustness studies
In real-world chemistry experiments, measurement frequency is often limited by instrumentation constraints and experimental costs, resulting in sparse observations, for example, sometimes as few as 10 time points over an entire reaction. To evaluate the robustness of our method under real-world conditions, we simulate increased sparsity by downsampling the concentration trajectories of the AOXID problem by factors of 2, 4, and 10, while keeping the total time span fixed. We then assess the accuracy of the CRNN pre-training under these conditions.
As shown in Table III, our method remains stable and yields physically plausible estimates even as data sparsity increases. The trajectory loss $\mathcal{L}_{y}$ and coefficient error $\mathcal{L}_{k}$ increase gradually, but the degradation in performance remains moderate. Notably, even with a 10-fold reduction in observations, the model achieves a reasonable rate coefficient error of 0.45. These results suggest that our approach is robust to sparse supervision and may be applicable to low-resolution experimental datasets.
Downsampling rate ODE fitting Rate coeff. estimat
w/o downsampling $8.9\times 10^{-5}$ 0.145
2 $5.3\times 10^{-4}$ 0.384
4 $6.4\times 10^{-4}$ 0.345
10 $2.0\times 10^{-3}$ 0.437
TABLE III: Robustness of CRNN pre-training under increasingly sparse observations. We downsample the data samples from the trajectories on the AOXID dataset, and report the MAE loss on ODE fitting and rate coefficients estimation in SPIN-ODE-s2.
V Conclusion
We propose a novel neural network-based framework for estimating chemical rate coefficients without the need for prior knowledge. To the best of our knowledge, this is the first approach capable of addressing stiff chemical ODEs under these conditions. This capability is particularly important in real-world chemical systems, such as atmospheric chemistry, where rate coefficients can vary exponentially and significantly influence reaction pathways. Our method employs a three-stage optimisation process: it begins by fitting stiff ODE trajectories, proceeds with explicit learning of the complete chemical system using a Chemical Reaction Neural Network (CRNN), and concludes with a refinement step that integrates CRNN with an ODE solver to enhance prediction accuracy. We validate our approach on both synthetic datasets and simplified real-world atmospheric chemical scenarios. The experimental results, both quantitative and qualitative, demonstrate that our method achieves state-of-the-art performance in ODE trajectory fitting and rate coefficient estimation. In future work, we aim to extend the approach to comprehensive chemical models and improve its robustness under more extreme stiffness conditions.
Acknowledgment
The authors wish to acknowledge CSC – IT Center for Science, Finland, for computational resources. W. P. thanks the Onni ja Hilja Tuovinen Foundation Grant for their generous support.